The present ages of a man and his son amount to 55 years.10 years ago,the man was 6 times as old as his son.find the present ages
Answer by Cromlix(4381)

The present ages of a man and his son amount to 55 years.
M + S = 55
M = 55 - S
10 years ago,the man was 6 times as old as his son.
M - 10 = 6(S - 10)
Replace 'M' on right with 55 - S
55 - S - 10 = 6S - 60
Collect like terms
-S -6S = -60 + 10 - 55
-7S = -105
S = 15
Present ages.
Son: 15yrs old
Man: 40yrs old
Hope this helps.
;-)